OpenStudy (anonymous):

Choose one of the factors of x6 + 1000 x2 x2 – 10 x4 – 10x2 + 100 x4 + 10x2 + 100

OpenStudy (anonymous):

factor as the sum of two cubes \[(x^2+10) (x^4-10 x^2+100)\]

OpenStudy (anonymous):

\(a^3+b^3=(a+b)(a^2-ab+b^2)\) with \(a=x^2, b=10\)

OpenStudy (e.cociuba):

One of the factors of x^6 + 1000: C) x^4 - 10x^2 + 100